OpenStudy (anonymous):

(chain rule and partial derivatives) If z = y + f(x^2 - y^2), where f is differentiable, show that y(del(z)/del(x)) + x(del(z)/del(y)) = x

OpenStudy (baru):

\[\frac{\partial z}{\partial y}=1+\frac{d f}{d(x^2-y^2)}(-2y) \\ \frac{\partial z}{\partial x}=\frac{d f}{d (x^2-y^2)}(2x)\]

OpenStudy (baru):

multiply first equation by x second equation by y add them up and see what you get

OpenStudy (anonymous):

OH MAN okay thanks a bunch, I see where I went astray now

OpenStudy (baru):

no prob :)